# SnapMoGen: Human Motion Generation from Expressive Texts

### v1

@@ 5.2 Component Analysis
We conduct comprehensive ablation experiments to evaluate the effects of various hyper-parameters and technical designs, as shown in Table 4. In row (A), we observe that compact pose representation leads to smaller VQ reconstruction error, while slightly underperforms for text-to-motion synthesis.
In terms of VQ configuration, we observe from rows (B) that while increasing codebook size consistently enhances VQ reconstruction and text-motion alignment (CLIP score), motion quality does not necessarily follow this trend (best FID at $|\mathcal{C}|=1024$). Rows (C) show that additional VQ layers effectively improve reconstruction, but more token hierarchies also introduce complexity for text-to-motion synthesis, with optimal results at 2 layers. In rows (D), we apply full-scale for all layers, yet despite achieving better VQ performance, inefficient token utilization
Figure 5: Validation loss curves on SnapMoGen. (B): Base. (M): Medium. (L): Large.
leads to suboptimal generation quality. Finally, incorporating self-attention layers in the encoder and decoder (row (E)) improves both VQ learning and motion synthesis performance.
We then study the components in the transformer. In row (F), we observe that the CLIP text encoder is clearly inadequate for handling the long, complex textual descriptions in SnapMoGen. In rows (G), we find that transformers with higher latent dimensions or more attention layers counterintuitively lead to worse motion generation quality. We can gain insights from Figure 5, which shows that transformer models with more parameters (Base: 18.3M, Medium: 36.6M, Large: 53.4M) are more prone to overfitting on the dataset.
@@ Appendix A Annotation Details
Our text annotation interface is presented in Fig. 6. We first visualize all motions using a 3D character to help annotators better understand the motion content. However, since the character may walk out of the camera view and inter-penetration artifacts sometimes occur, we also display the motions using stick-figure representations that remain centered in the camera view. These two visualizations are synchronized and presented simultaneously to annotators. Annotators can also flag low-quality motions during the annotation process.
@@ B.2 Evaluation Model
Our evaluation model accounts for both motion fidelity and text-motion alignment. We adopt the TMR framework, as shown in Figure 7. This framework comprises three network components: a motion encoder that encodes motion sequences into global vectors, a text encoder that encodes text sequences into global vectors, and a motion decoder that reconstructs motions from either motion or text vectors. All three networks are 6-layer transformers with a latent dimension of 256, 4 attention heads, and a feedforward hidden size of 1024. The T5-base model first extracts word-level features from texts. For motions, we use only the essential 148-dimensional root motion and local rotational features. All encoders output Gaussian distribution parameters (mean and log-variance), from which vectors are sampled. We append two temporal timesteps at the end of the sequence input for outputing these vectors.
@@
where $\mathcal{L}_{\textbf{rec}}$ measures the motion reconstruction given text or motion input (via a smooth L1 loss). A KL-divergence loss $\mathcal{L}_{\text{KL}}$ regularizes each embedding distribution to be close to a unitary Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, and also encourage these two distributions to be close to each other. $\mathcal{L}_{\text{E}}$ enforces both mean vectors to be similar to each other. Finally, a InfoNCE oord2018representation loss is used for constrastive learning of motion-text batches with batch size of 64. We set $\lambda_{\text{E}}$, $\lambda_{\text{KL}}$, and $\lambda_{\text{NCE}}$ to $51e-$ 5, $51e-$ 5, and $0.1$. For more model details, we recommend to read the original TMR work petrovich2023tmr .
@@ Appendix C LLM-based Prompt Augmentation
During training, we enhance data diversity by employing an LLM to rewrite human-provided annotations, generating paraphrased versions with varied linguistic structures while preserving core semantics. This approach ensures each motion sequence is associated with multiple textual descriptions, improving model robustness. The prompt instructions for ChatGPT are provided in Tab. 5.
@@
During inference, the LLM rewrites each input prompt into a richly detailed description, incorporating explicit motion cues such as body posture, timing, and stylistic elements. This expanded form more effectively guides motion generation models. The instructions for ChatGPT are provided in Tab. 6.

### v2

@@ 5.2 Component Analysis
We perform comprehensive ablation experiments to evaluate the effects of various hyper-parameters and technical designs, as shown in Tab.˜4 and  Tab.˜5. In  Tab.˜4 (A), we observe that compact pose representation leads to a smaller VQ reconstruction error, while it slightly underperforms for text-to-motion synthesis.
In terms of VQ configuration, we observe from Tab.˜4 (B) that while increasing codebook size consistently enhances VQ reconstruction and text-motion alignment (CLIP score), motion quality does not necessarily follow this trend (best FID at $|\mathcal{C}|=1024$). Tab.˜4 (C) show that additional VQ layers effectively improve reconstruction, but more token hierarchies also introduce complexity for text-to-motion synthesis, with optimal results at 2 layers. In Tab.˜4 (D), we apply full-scale for all layers, yet despite achieving better VQ performance, inefficient token utilization leads to suboptimal generation quality. Finally, incorporating self-attention layers in the encoder and decoder (Tab.˜4 (E)) improves both VQ learning and motion synthesis performance.
Figure 5: Decoding progress over iterations for different token scales.
Figure 6: Validation loss curves of different model sizes on SnapMoGen.
We then examine the effects of text augmentation and text-to-motion transformer design in Tab.˜5. In Tab.˜5 (A), caption augmentation clearly improves model performance across all evaluation metrics. In Tab.˜5 (B), we observe that the CLIP text encoder is inadequate for handling the long and complex textual descriptions in SnapMoGen. From Tab.˜5 and Fig.˜6, we further find that cross-attention conditioning is less prone to overfitting and leads to higher motion quality and better text–motion alignment. Meanwhile, Tab.˜5 (D) show that transformers with higher latent dimensions or more attention layers counterintuitively degrade motion generation quality. Figure˜6 provides additional insight, indicating that larger transformer models (Base: 18.3M, Medium: 36.6M, Large: 53.4M) tend to overfit the dataset more severely.
How are tokens at different scales modeled?
To investigate this question, we analyze which tokens are “favored” by MoMask++. During iterative inference, MoMask++ generates a complete motion sequence by selectively retaining and re-masking tokens at each step, allowing us to track which tokens the model prioritizes over time. We conduct an experiment using four token scales (from coarse to fine, with 10, 20, 40, and 80 tokens, totaling 150 tokens per sequence) and a 10-iteration inference process over 32 text prompts, recording the token completion ratio at each scale. The results, shown in Fig.˜6, reveal that the model naturally prioritizes coarse-scale tokens (Scale 1) in the early stages of generation and progressively shifts its focus toward finer scales. This behavior demonstrates a “global-to-local” generation strategy, indicating that the attention mechanism effectively captures and prioritizes information based on semantic importance (coarse-to-fine).
@@ Appendix A Annotation Details
Our text annotation interface is presented in Fig.˜7. We first visualize all motions using a 3D character to help annotators better understand the motion content. However, since the character may walk out of the camera view and inter-penetration artifacts sometimes occur, we also display the motions using stick-figure representations that remain centered in the camera view. These two visualizations are synchronized and presented simultaneously to annotators. Annotators can also flag low-quality motions during the annotation process.
@@ B.2 Evaluation Model
Our evaluation model accounts for both motion fidelity and text-motion alignment. We adopt the TMR framework, as shown in Figure˜8. This framework comprises three network components: a motion encoder that encodes motion sequences into global vectors, a text encoder that encodes text sequences into global vectors, and a motion decoder that reconstructs motions from either motion or text vectors. All three networks are 6-layer transformers with a latent dimension of 256, 4 attention heads, and a feedforward hidden size of 1024. The T5-base model first extracts word-level features from texts. For motions, we use only the essential 148-dimensional root motion and local rotational features. All encoders output Gaussian distribution parameters (mean and log-variance), from which vectors are sampled. We append two temporal timesteps at the end of the sequence input for outputing these vectors.
@@
where $\mathcal{L}_{\textbf{rec}}$ measures the motion reconstruction given text or motion input (via a smooth L1 loss). A KL-divergence loss $\mathcal{L}_{\text{KL}}$ regularizes each embedding distribution to be close to a unitary Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, and also encourage these two distributions to be close to each other. $\mathcal{L}_{\text{E}}$ enforces both mean vectors to be similar to each other. Finally, a InfoNCE oord2018representation loss is used for constrastive learning of motion-text batches with batch size of 64. We set $\lambda_{\text{E}}$, $\lambda_{\text{KL}}$, and $\lambda_{\text{NCE}}$ to 1e51e-5, 1e51e-5, and $0.1$. For more model details, we recommend to read the original TMR work petrovich2023tmr .
@@ Appendix C LLM-based Prompt Augmentation
During training, we enhance data diversity by employing an LLM to rewrite human-provided annotations, generating paraphrased versions with varied linguistic structures while preserving core semantics. This approach ensures each motion sequence is associated with multiple textual descriptions, improving model robustness. The prompt instructions for ChatGPT are provided in Tab.˜6.
@@
During inference, the LLM rewrites each input prompt into a richly detailed description, incorporating explicit motion cues such as body posture, timing, and stylistic elements. This expanded form more effectively guides motion generation models. The instructions for ChatGPT are provided in Tab.˜7.